# Machine Unlearning: A Comprehensive Survey

### v2

@@ 10. Summary
The survey aims to offer a comprehensive and systematic overview of machine unlearning techniques. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of corresponding unlearning methods. Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning to be an important part of the studies. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive picture of existing work and listing the challenges and open problems of different scenarios of machine unlearning.

### v3

@@ 10. Summary
The survey aims to offer a comprehensive and systematic overview of machine unlearning techniques. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we briefly categorize recent studies into exact and approximate unlearning and introduce the technical details of the corresponding unlearning methods. Moreover, we noticed some new unlearning scenarios, such as graph unlearning, federated unlearning, diffusion model unlearning, and LLMs unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning to be an important part of the studies. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive picture of existing work and listing the challenges and open problems of different scenarios of machine unlearning.